# Regularization of the ensemble Kalman filter using a non-parametric, non-stationary spatial model

### v3

@@ 3.2 Sample band variances
Consider the application of ${\cal H}_{j}$ to $\xi(x)$. As shown in H, if $L_{{\cal H}_{j}}\ll\Lambda$, (where $L_{{\cal H}_{j}}$ is the length scale of the $j$th filter’s impulse response function and $\Lambda$ the non-stationarity length scale introduced in G), then the spectral-space filtering can be approximately performed as if the random field $\xi(x)$ were stationary:
$\varphi_{j}(x)\approx\sum^{\ell_{\rm max}}_{\ell=0}\sum^{\ell}_{m=-\ell}H_{j}(\ell)\,\sigma_{\ell}(x)\,\widetilde{\alpha}_{\ell m}\,Y_{\ell m}(x).$ (8)
This quation implies that there is an approximate relation between the local spectrum at the spatial point $x$ and variances of the processes $\varphi_{j}(x)$ at the same point:
$v_{j}(x)={\mathsf{Var}}\,(\varphi_{j}(x){\mbox{$\,|\,$}}\sigma)\approx\sum_{\ell=0}^{\ell_{\rm max}}\frac{2\ell+1}{4\pi}H_{j}^{2}(\ell)\,\sigma_{\ell}^{2}(x),$ (9)
where conditioning on $\sigma$ is meant in the sense that all functions $\sigma_{\ell}(\cdot)$ are fixed. According to Eq. (67), the mean magnitude of the approximation error $\delta^{\rm nst}_{j}$ in Eq. (9), which arises due to non-stationarity of $\xi(x)$, is as small as $\left({L_{{\cal H}_{j}}}/{\Lambda}\right)^{2}$.
As the band variances $v_{j}(x)$ are readily estimated from the ensemble by their sample counterparts, $d_{j}(x)$, and $H_{j}(\ell)$ are known, we can use Eq. (9) to formulate a linear inverse problem with respect to $f_{\ell}(x)=\sigma_{\ell}^{2}(x)$ formulated for each spatial grid point $x$ independently:
${\bf d}(x)=\boldsymbol{\Omega}\,{\bf f}(x)+\boldsymbol{\delta}^{\rm smp}(x)+\boldsymbol{\delta}^{\rm nst}(x),$ (10)
where ${\bf d}(x)=\{d_{j}(x)\}_{j=1}^{J}$ is the vector of (observed) sample band variances, $\boldsymbol{\Omega}$ is an $J$ by $\ell_{\rm max}+1$ matrix, $\boldsymbol{\delta}^{\rm smp}(x)$ is the vector of sampling uncertainties in ${\bf d}(x)$, and $\boldsymbol{\delta}^{\rm nst}(x)=\{\delta^{\rm nst}_{j}(x)\}_{j=1}^{J}$ is the non-stationarity-error vector. Note that Eqs. (9) and (10) imply that $d_{j}(x)$ are nothing other than the aggregated local spectra c $f_{\ell}(x)=\sigma_{\ell}^{2}(x)$ contaminated with sampling noise $\delta_{j}^{\rm smp}(x)$ and the approximation error $\delta^{\rm nst}_{j}$. Informally, for a given $x$, a disaggregation of $\boldsymbol{\sigma}(x)=\{\sigma_{\ell}(x)\}_{\ell=0}^{\ell_{\rm max}}$ from ${\bf d}(x)$ should be possible because $\sigma_{\ell}(x)$ is a smooth function of wavenumber.
In I, we show that, with a theoretically convenient spectral smoothing constraint, a consistent estimator can be built by solving the linear inverse problem Eq. (10). To achieve consistency, the widths of the bandpass filters’ transfer functions are to be selected depending on the ensemble size: the bigger the ensemble, the broader bandpass filters’ spectral transfer functions can and should be taken. Broader transfer functions imply more localized filters’ impulse response functions and therefore less approximation error $\delta^{\rm nst}_{j}$ in $d_{j}(x)$ due to the spatial non-stationarity, resulting in more accurately estimated local spectra.
In the next subsection, a practically relevant estimator of $\sigma_{\ell}(x)$ from ${\bf d}(x)$ is developed as a proxy to a Bayes estimator.
@@ 3.3 Neural Bayes estimator of local spectrum from sample band variances
We recall that the data ${\bf d}(x)$ are sample variances of the fields $\varphi_{j}^{(\mu)}(x)$, which are the bandpass filtered ensemble members $\xi^{(\mu)}(x)$ generated by our Locally Stationary Convolution Model with the shared $\boldsymbol{\sigma}(x)$. We denote this sequence of transformations as ${\bf d}(x)=\boldsymbol{\omega}(\boldsymbol{\sigma}(\cdot),x)$ where $\boldsymbol{\omega}$ is the data model (a known random operator). The data model induces a probability distribution of the data ${\bf d}(x)$ at any point $x$ given the field of model’s parameters $\boldsymbol{\sigma}(\cdot)$, i.e., $p({\bf d}(x){\mbox{$\,|\,$}}\boldsymbol{\sigma}(\cdot))$, the likelihood of the data222 In the context of sequential filtering, $p(\boldsymbol{\sigma}(\cdot))$ is actually a hyperprior, meaning that $\boldsymbol{\sigma}(\cdot)$ defines, according to our Locally Stationary Convolution Model, the prior Gaussian distribution of the unknown system state before current observations on the state are assimilated.. Suppose, also, that we have a prior distribution, $p(\boldsymbol{\sigma}(\cdot))$, over the random spectral functions $\sigma_{\ell}(x)$. The prior distribution encodes the four constraints of our spatial model, see section 2.8. Having both the prior and the likelihood, we aim to build a proxy to the Bayes point estimator.
We postulate that the estimator is of the form
$\widehat{\boldsymbol{\sigma}}(x)=\Phi({\bf d}(x),\boldsymbol{\psi}),$ (11)
@@
where $\Phi$ is a known function of its arguments and $\boldsymbol{\psi}$ is the unknown vector of the estimator’s parameters. To optimize the estimator, that is, to find the best $\boldsymbol{\psi}$, we introduce a loss function chosen to be quadratic:
${\cal L}(\widehat{\boldsymbol{\sigma}}(x),\,\boldsymbol{\sigma}(x))=\sum\frac{2\ell+1}{4\pi}\left(\widehat{\sigma_{\ell}}(x)-\sigma_{\ell}(x)\right)^{2}.$ (12)
@@
Then, we minimize the Bayes risk, that is, the expectation of the loss function with respect to the joint distribution of the parameter $\boldsymbol{\sigma}(\cdot)$ and the data ${\bf d}(x)=\boldsymbol{\omega}(\boldsymbol{\sigma}(\cdot),x)$:
$\widehat{\boldsymbol{\psi}}=\operatorname*{argmin}\operatorname{\mathbb{E}}\,{\cal L}\left(\Phi({\bf d}(x),\boldsymbol{\psi}),\,\boldsymbol{\sigma}(x)\right),$ (13)
@@
getting the mean-square optimal estimator’s parameters $\boldsymbol{\psi}$. This is the classical Bayes estimator (e.g., Lehmann and Casella, 2006, Chapter 4).
@@
As tackling non-Gaussian densities in high dimensions is rarely feasible, we rely on a Monte-Carlo approximation. In doing so, we assume that, first, there is a large sample $\{\boldsymbol{\sigma}^{i}(\cdot)\}_{i=1}^{I}$ ($I$ is its size) drawn from the prior distribution $p(\boldsymbol{\sigma}(\cdot))$. Second, we note that the data model $\boldsymbol{\omega}(\boldsymbol{\sigma}(\cdot),x)$ is generative: given $\boldsymbol{\sigma}^{i}(\cdot)$, it can simulate the data ${\bf d}^{ij}(x_{r})$, which are $J$ sample band variances at $R$ selected grid points $x_{r}$, where $r=1,2,\dots,R$. Since, by construction, $\boldsymbol{\sigma}^{i}(\cdot)\sim p(\boldsymbol{\sigma}(\cdot))$ and ${\bf d}^{ij}(x_{r})\sim p({\bf d}(x_{r}){\mbox{$\,|\,$}}\boldsymbol{\sigma}(\cdot))$, we can approximate the expectation in Eq. (13)) by its Monte-Carlo estimate, getting
$\frac{1}{n_{\rm train}}\sum_{ijr}{\cal L}\left(\Phi({\bf d}^{ij}(x_{r}),\boldsymbol{\psi}),\;\boldsymbol{\sigma}^{i}(x_{r})\right)\to\min_{\boldsymbol{\psi}},$ (14)
@@
where $n_{\rm train}=IJR$ is the size of the training sample. In this study, we choose a neural network as the parametric estimator $\Phi({\bf d}(x),\boldsymbol{\psi})$ with $\boldsymbol{\psi}$ being the collection of the net’s weights and biases. If (i) the training sample is large and representative, (ii) the neural network is expressive enough, and (ii) the neural net is capable of finding a good proxy to the global minimum in Eq. (14), then the resulting estimator is a good approximation to the Bayes estimator. In this way, we obtain a neural Bayes estimator (Sainsbury-Dale et al., 2024). Technical details of the neural net are given in K.
@@
The above neural-network based disaggregation of the local spectrum from the set of sample band variances makes the technique applicable to high-dimensional problems because the neural net is fast and the computations can be performed at different grid points in parallel.
4 Application in the sequential filter
The goal of this study is a technique aimed to regularize (by using a spatial random field model) the prior covariance matrix in sequential ensemble filtering. Testing and validation of such a technique can well be performed in the linear setting, avoiding unnecessary complications related to nonlinearities of observation and forecast models. For this reason, in the sequel we confine ourselves to the linear filtering setup. A reviews optimal and sub-optimal, ensemble and non-ensemble, time-discrete linear filters that are used in the numerical experiments presented below.

### v5

@@ 4.2 Approximate likelihood of the filtered signals
Consider the application of the filter ${\cal H}_{j}$ to $\xi(x)$. As shown in F, if $L_{{\cal H}_{j}}\ll\Lambda$, (where $L_{{\cal H}_{j}}$ is the length scale of the $j$th bandpass filter’s impulse response function and $\Lambda$ the process’s non-stationarity length scale introduced in section 3.6), then the spectral-space filtering can be approximately performed as if the random field $\xi(x)$ were stationary:
$\varphi_{j}(x)=\sum^{\ell_{\rm max}}_{\ell=0}\sum^{\ell}_{m=-\ell}H_{j}(\ell)\,\sigma_{\ell}(x)\,\widetilde{\alpha}_{\ell m}\,Y_{\ell m}(x)+\delta\varphi_{j}(x),$ (15)
where $\delta\varphi_{j}(x)$ is the discrepancy due to the non-stationarity, which has mean zero and variance $\operatorname{\mathbb{E}}|\delta\varphi_{j}|^{2}=O\left(L_{{\cal H}_{j}}^{2}/\Lambda^{2}\right)$ as $L_{{\cal H}_{j}}/\Lambda\to 0$, see Eq. (69). The key point here is that, up to the small discrepancy, the distribution of $\varphi_{j}(x)$ is determined by the spectral functions $\sigma_{\ell}(x)$ evaluated at the same spatial point $x$ only.
Specifically, given $\boldsymbol{\sigma}(x)=\{\sigma_{\ell}(x)\}_{\ell=0}^{\ell=\ell{\rm max}}$, the random vector $\boldsymbol{\varphi}(x)=\{\varphi_{j}(x)\}_{j=1}^{J}$ is multivariate Gaussian with mean zero and covariances $\operatorname{\mathbb{E}}\varphi_{j}(x)\,\varphi_{j^{\prime}}(x)=\sum^{\ell_{\rm max}}_{\ell=0}\frac{2\ell+1}{4\pi}H_{j}(\ell)H_{j^{\prime}}(\ell)\,\sigma_{\ell}^{2}(x)+O(L_{{\cal H}_{j}}/\Lambda)+O(L_{{\cal H}_{j^{\prime}}}/\Lambda)$. If the small $O(\cdot)$ contributions are neglected, this fully determines its likelihood $p(\boldsymbol{\varphi}(x){\mbox{$\,|\,$}}\boldsymbol{\sigma}(x))$ for each grid point $x$.
For computational reasons, instead of relying on the exact likelihood of $\boldsymbol{\varphi}(x)$, we neglect cross-correlations between different $\varphi_{j}(x)$ at the same grid point $x$. Besides, considering the likelihood grid point by grid point, we, effectively, neglect cross-correlations between different grid points as well. As a result, we end up with the approximate Gaussian composite likelihood defined for each spatial grid point separately. For the whole sample of $K$ ensemble members the composite log-likelihood reads (up to a constant)
$c\ell[\boldsymbol{\sigma}(x)]=-\frac{K}{2}\sum_{j=1}^{J}\left(\ln v_{j}(x)+\frac{d_{j}(x)}{v_{j}(x)}\right),$ (16)
where $v_{j}(x)={\mathsf{Var}}\,(\varphi_{j}(x){\mbox{$\,|\,$}}\boldsymbol{\sigma}(x))$ are the band variances and $d_{j}(x)$ are their sample counterparts, the sample band variances. From Eq. (15),
$v_{j}(x)=\utilde{v}_{j}(x)+\delta^{\rm nsta}_{j}(x),$ (17)
where
$\utilde{v}_{j}(x)=\sum_{\ell=0}^{\ell_{\rm max}}\frac{2\ell+1}{4\pi}H_{j}^{2}(\ell)\,\sigma_{\ell}^{2}(x)$ (18)
is the approximate band variance and $\delta^{\rm nsta}_{j}(x)=O({L_{{\cal H}_{j}}}/{\Lambda})$ is the approximation error.
According to Eqs. (17) and (18), the band variances $v_{j}(x)$ are, approximately, the aggregated local spectra $f_{\ell}(x)=\sigma_{\ell}^{2}(x)$. Jointly, $\{v_{j}(x)\}_{j=1}^{J}$ bear all the information on the (smooth) local spectrum we wish to infer from the sample (up to the small discrepancy $\delta^{\rm nsta}_{j}$). On the other hand, the sample band variances $d_{j}(x)$ summarize information on the local spectrum contained in the raw sample $\{\xi^{(k)}(x)\}_{k=1}^{K}$ (up to the sampling error: $d_{j}(x)=v_{j}+\delta^{\rm samp}_{j}$). Note that $v_{j}(x)$ and $d_{j}(x)$ are the locally stationary counterparts of the stationary spectrum $f_{\ell}$ and periodogram $\widehat{v}_{\ell}$, respectively, see section 4.1.
4.3 A two-stage estimator
Equations (16)–(18) suggest the following two-stage estimator applied point-wise, for each spatial grid point $x$ independently:
  1. 1.
    Compute the sample band variances $d_{j}(x)$:
    1. (a)
      Bandpass filter $K$ ensemble perturbations, $\xi^{(k)}(\cdot)$, $k=1,2\dots,K$. The result is the $JK$ filtered fields $\varphi_{j}^{(k)}(\cdot)$, $j=1,2\dots,J$.
    2. (b)
      At each spatial grid point $x$, compute sample variances of $\varphi_{j}^{(k)}(x)$, that is, the sample band variances $d_{j}(x)$.
  2. 2.
    Recover $\boldsymbol{\sigma}(x)$ from the set of sample band variances ${\bf d}(x)=\{d_{j}(x)\}_{j=1}^{J}$. We consider two ways of doing so.
    • The first approach relies on linearity of Eq. (18) with respect to the local spectrum $f_{\ell}(x)=\sigma_{\ell}^{2}(x)$ and neglect of $\delta^{\rm nsta}_{j}(x)$. We rewrite Eqs. (17) and (18) as
      ${\bf v}(x)\approx\boldsymbol{\Omega}\,{\bf f}(x),$ (19)
      where ${\bf v}(x)=\{v_{j}(x)\}_{j=1}^{J}$, ${\bf f}(x)=\{f_{\ell}(x)\}^{\ell_{\rm max}}_{\ell=0}$, and $\boldsymbol{\Omega}$ is the $J$ by $\ell_{\rm max}+1$ matrix with the known entries $(\boldsymbol{\Omega})_{j\ell}=\frac{2\ell+1}{4\pi}H_{j}^{2}(\ell)$.
      Equations (16) and (19) imply that the composite likelihood $c\ell[\boldsymbol{\sigma}(x)]$ is approximately maximized at any vector ${\bf f}(x)$ that satisfies
      ${\bf d}(x)=\boldsymbol{\Omega}\,{\bf f}(x).$ (20)
      This estimating equation is underdetermined and thus needs to be regularized. In section 4.4, we do so by imposing a deterministic smoothness constraint and show that the error of the resulting solution vanishes in the large sample limit so that the whole estimator is consistent.
    • The second approach is to rely on the available unapproximated nonlinear data generation (or ‘forward’) model $\boldsymbol{\omega}$ such that
      ${\bf d}(x)=\boldsymbol{\omega}(\boldsymbol{\sigma}(\cdot),x).$ (21)
      In section 4.5, we build an optimized neural-network based mapping ${\bf d}(x)\mapsto\boldsymbol{\sigma}(x)$ that approximates the classical Bayes estimator.
    In the next two subsections, we elaborate on the two implementations of stage 2, one based on Eq. (20) and the other on Eq. (21).
4.4 Consistency of the linear estimator
To solve the underdetermined Eq. (20) we employ the deterministic regularization technique known as ‘spectral cut-off’, (e.g. Cavalier, 2011). Specifically, we impose a smoothness constraint that requires the local spectrum to be band-limited as a function of log-wavenumber, see below.
To formulate the smoothness constraint, we switch from wavenumbers $\ell$ to log-wavenumbers $\vartheta(\ell)=\pi\log(\ell+1)/\log(\ell_{\rm max}+1)$. We set $\vartheta(-\ell)=\vartheta(\ell)$ so that $\vartheta$ ranges from $-\pi$ to $\pi$. Then, we let $F(\vartheta(\ell))=f_{\ell}$. We also assume that the spatial resolution is large enough and the spectra are smooth enough to treat $F(\vartheta)$ as a function of continuous argument on the unit circle.
Definition 4.1.
We say that the spectrum $f_{\ell}$ is smooth with the bandwidth $M$ if $F(\vartheta)$ is a band-limited function:
$F(\vartheta)=\sum_{m=-M}^{M}\widetilde{F}_{m}\,\mathrm{e}^{\mathsf{i}m\vartheta},$ (22)
where $\widetilde{F}_{m}$ is the the Fourier transform of $F(\vartheta)$ and $M$ is a positive integer.
The smoothness constraint in this section is the assumption that the spectrum is smooth with the bandwidth $M$ (the lower $M$, the smoother $F(\cdot)$). Since $F(\vartheta)$ is an even real-valued function, it follows from Eq. (22) that the spectrum becomes parameterized with the vector, ${\bf p}$, of $M+1$ real-valued parameters: ${\bf f}=\boldsymbol{\Psi}{\bf p}$ with the known $(\ell_{\rm max}+1)\times(M+1)$ matrix $\boldsymbol{\Psi}$. Therefore, we can select $J=M+1$ bandpass filters so that the number of unknowns becomes equal to the number of data (the sample band variances ${\bf d}$). Then, Eq. (20) reduces to a system of linear algebraic equations:
${\bf d}=\boldsymbol{\Omega}\boldsymbol{\Psi}{\bf p}.$ (23)
The solution to this equation is subject to both sampling error $\boldsymbol{\delta}^{\rm samp}(x)=\{\delta^{\rm samp}_{j}(x)\}_{j=1}^{J}$ and non-stationarity error $\boldsymbol{\delta}^{\rm nsta}(x)=\{\delta^{\rm nsta}_{j}(x)\}_{j=1}^{J}$. In G we prove that their impact on the estimated local spectrum results in the estimation error that can be made arbitrarily small by increasing the sample size and appropriately increasing spectral widths of the filters. More specifically, we prove the following
Proposition 4.1.
Let the local spectrum be smooth in terms of Definition 4.1 with the bandwidth $M$. Then, for every sample size $K$, there exist $M+1$ bandpass filters such that the solution to Eq. (23) yields vanishing mean absolute error in the local spectrum as $K\to\infty$. The spectral widths of the filters (in terms of log-wavenumbers) grow with the increasing sample size $K$ as $K^{\gamma}$ for some $\gamma>0$.
Due to Markov’s inequality, Proposition 4.1 entails that the estimator is consistent.
Two points are worth noting. First, while proving consistency we found that the widths of the bandpass filters’ transfer functions are to be selected depending on the ensemble size: the bigger the ensemble, the broader bandpass filters’ spectral transfer functions can and should be taken. Broader transfer functions imply more localized filters’ impulse response functions and therefore less non-stationarity error $\delta^{\rm nsta}_{j}$ in $d_{j}(x)$. This allows the estimator to extract increasingly detailed spatially non-stationary structures, resulting in more accurately estimated local spectra. Second, this conclusion has important consequences not just for the theoretical proof of consistency but also for a more practical implementation of stage 2 of the estimator (presented in the next subsection), where such a conclusion could not be derived analytically.
@@ 4.5 Neural Bayes estimator of local spectrum from sample band variances
We recall that the second implementation of stage 2 of the estimator, see section 4.3, abandons the linear approximation made in the previous subsection and relies on Eq. (21. This makes the ‘forward model’ more accurate and reduces the noise level — but at the expense of rendering the problem nonlinear and high-dimensional. Indeed, the exact expression for the filtered process $\varphi_{j}(x)$ is
$\varphi_{j}(x)=\int_{D}h_{j}(\rho(x,y))\xi(y)\,{\rm d}y=\sum^{\ell_{\rm max}}_{\ell=0}\sum^{\ell}_{m=-\ell}\widetilde{\alpha}_{\ell m}\int_{D}h_{j}(\rho(x,y))\,\sigma_{\ell}(y)\,Y_{\ell m}(y)\,{\rm d}y,$ (24)
where Eq. (10) was used. From Eq. (24), it follows that the respective band variance $v_{j}(x)=\operatorname{\mathbb{E}}(|\varphi_{j}(x)|^{2}{\mbox{$\,|\,$}}\boldsymbol{\sigma}(.))$ depends not only on $\sigma_{\ell}^{2}(x)$ (i.e., evaluated at the same point $x$ at which $\varphi_{j}$ is evaluated), but also on the products $\sigma_{\ell}(y)\,\sigma_{\ell}(y^{\prime})$ for all pairs of spatial points $(y,y^{\prime})$. This justifies writing the data model as ${\bf d}(x)=\boldsymbol{\omega}(\boldsymbol{\sigma}(\cdot),x)$, as in Eq. (21). In high dimensions, this also implies that the likelihood of ${\bf d}(x)$ is numerically intractable. Therefore we build an estimator of $\boldsymbol{\sigma}(x)$ that does not require evaluations of the likelihood function.
We start with postulating that the estimator is of the form
$\widehat{\boldsymbol{\sigma}}(x)=\boldsymbol{\Phi}({\bf d}(x),\boldsymbol{\psi}),$ (25)
@@
where $\boldsymbol{\Phi}=(\Phi_{0},\Phi_{1},\dots,\Phi_{\ell_{\rm max}})$ is a known vector function of its arguments and $\boldsymbol{\psi}$ is the unknown vector of the estimator’s parameters.
We place a prior distribution, $p(\boldsymbol{\sigma}(\cdot))$, over the unknown spectral functions $\sigma_{\ell}(\cdot)$. Note that $p(\boldsymbol{\sigma}(\cdot))$ is a hyperprior in the context of ensemble filtering. The hyperprior $p(\boldsymbol{\sigma}(\cdot))$ encodes the four constraints of our spatial model, see section 3.8.
To optimize the estimator, that is, to find the best $\boldsymbol{\psi}$, we introduce a loss function chosen to be quadratic:
${\cal L}[\widehat{\boldsymbol{\sigma}}(\cdot),\,\boldsymbol{\sigma}(\cdot)]=\frac{1}{|D|}\int_{D}\sum_{\ell=0}^{\ell_{\rm max}}\frac{2\ell+1}{4\pi}\left(\widehat{\sigma_{\ell}}(x)-\sigma_{\ell}(x)\right)^{2}\,{\rm d}x$ (26)
@@
(where $|D|$ is the area of the analysis domain $D$). Aiming at the Bayes point estimator (e.g. Lehmann and Casella, 2006, Chapter 4), we minimize the Bayes risk, that is, the expectation of the loss function with respect to the joint distribution of the parameter $\boldsymbol{\sigma}(\cdot)$ and the data ${\bf d}(\cdot)$:
$\widehat{\boldsymbol{\psi}}=\operatorname*{argmin}_{\boldsymbol{\psi}}\operatorname{\mathbb{E}}\,{\cal L}[\boldsymbol{\Phi}({\bf d}(\cdot),\boldsymbol{\psi}),\;\boldsymbol{\sigma}(\cdot)],$ (27)
@@
getting the mean-square optimal estimator’s parameters $\boldsymbol{\psi}$ and thus the optimal estimator defined by Eq. (25).
@@
As the likelihood function is numerically intractable, we rely on a Monte-Carlo approximation to solve the optimization problem Eq. (27). In doing so, we assume that, first, there is a sample $\{\boldsymbol{\sigma}^{i}(\cdot)\}_{i=1}^{I}$ ($I$ is its size) drawn from the hyperprior distribution $p(\boldsymbol{\sigma}(\cdot))$. Second, we exploit the fact that the data model $\boldsymbol{\omega}(\boldsymbol{\sigma}(\cdot),x)$ is generative: given $\boldsymbol{\sigma}^{i}(\cdot)$, it can simulate the data ${\bf d}^{i}(x_{r})$, that is, $J$ sample band variances at $R$ randomly and uniformly selected grid points $x_{r}$, where $r=1,2,\dots,R$. Since, by construction, $\boldsymbol{\sigma}^{i}(\cdot)\sim p(\boldsymbol{\sigma}(\cdot))$ and ${\bf d}^{i}(\cdot)\sim p({\bf d}(\cdot){\mbox{$\,|\,$}}\boldsymbol{\sigma}^{i}(\cdot))$, we can approximate the expectation in Eq. (27)) and the integral in Eq. (26) using Monte-Carlo:
$\frac{1}{n_{\rm train}}\sum_{i=1}^{I}\sum_{r=1}^{R}\sum_{\ell=0}^{\ell_{\rm max}}\frac{2\ell+1}{4\pi}\left(\Phi_{\ell}({\bf d}^{i}(x_{r}),\boldsymbol{\psi})-\sigma^{i}_{\ell}(x_{r})\right)^{2},$ (28)
where $n_{\rm train}=IR$ is the size of the training sample.
@@
In this study, we choose a neural network as the parametric estimator $\boldsymbol{\Phi}({\bf d}(x),\boldsymbol{\psi})$ with $\boldsymbol{\psi}$ being the collection of the net’s weights and biases. If (i) the training sample $\{\boldsymbol{\sigma}^{i}(\cdot)\}_{i=1}^{I}$ is large and representative of the hyperprior $p(\boldsymbol{\sigma})$, (ii) the neural network is expressive enough, and (iii) the network optimizer is capable of finding a good proxy to the global minimum in Eq. (28), then the resulting estimator is a good approximation to the Bayes estimator. In this way, we obtain a neural Bayes estimator (Sainsbury-Dale et al., 2024). Technical details on the neural net are given in I.
@@
The above neural-Bayes based disaggregation of the local spectrum from the set of sample band variances makes the technique applicable to high-dimensional problems. This is because (i) the estimator is local in space, that is, for each spatial point $x$, it maps ${\bf d}(x)$ to $\widehat{\boldsymbol{\sigma}}(x)$, (ii) the resulting inference is amortized: no retraining is needed when new data arrive (unless there are changes in the hyperprior $p(\boldsymbol{\sigma}(\cdot))$ or in the ensemble size, or in the bandpass filters), (iii) the computations can be performed at different grid points in parallel.
It is worth reiterating that the estimator $\widehat{\boldsymbol{\sigma}}(x)=\boldsymbol{\Phi}({\bf d}(x),\boldsymbol{\psi})$ is trained only once, that is, its parameters $\boldsymbol{\psi}$ are optimized during the training and then kept fixed. The application of the estimator to fresh data ${\bf d}(x)$ requires no additional optimization. The neural network just transforms its input ${\bf d}(x)$ to the estimate $\widehat{\boldsymbol{\sigma}}(x)$.
At the final stage of the estimator, we substitute $\widehat{\sigma}_{\ell}(x)$ into Eq. (6), getting the estimated spatially varying convolution kernel $\widehat{u}(x,\rho)$, which, in turn, is used to compute the $\widehat{\bf W}$ matrix, the key ingredient of the analysis algorithm, see section 5. The resulting spatial covariances, being determined by both the hyperprior (‘climatological’) distribution of the local spectra and the current ensemble, are non-stationary in time as well as in space. The uncertainty of the estimate is not taken into account in this study.
4.6 Summary of the estimator’s workflow
The practical estimation procedure consists of the following steps.
  • Ensemble perturbations are bandpass filtered.
  • Sample variances of the filtered fields are computed at each grid point.
  • Local spectra are inferred from the sample band variances using a neural network grid point by grid point.
  • The spatially varying convolution kernel is computed from the estimated local spectra.
5 LSEF: A Locally Stationary Ensemble Filter